# Unleashing the Temporal-Spatial Reasoning Capacity of GPT for Training-Free Audio and Language Referenced Video Object Segmentation

### v1

@@ Abstract
In this paper, we propose an Audio-Language-Referenced SAM 2 (AL-Ref-SAM 2) pipeline to explore the training-free paradigm for audio and language-referenced video object segmentation, namely AVS and RVOS tasks. The intuitive solution leverages GroundingDINO to identify the target object from a single frame and SAM 2 to segment the identified object throughout the video, which is less robust to spatiotemporal variations due to a lack of video context exploration. Thus, in our AL-Ref-SAM 2 pipeline, we propose a novel GPT-assisted Pivot Selection (GPT-PS) module to instruct GPT-4 to perform two-step temporal-spatial reasoning for sequentially selecting pivot frames and pivot boxes, thereby providing SAM 2 with a high-quality initial object prompt. Within GPT-PS, two task-specific Chain-of-Thought prompts are designed to unleash GPT’s temporal-spatial reasoning capacity by guiding GPT to make selections based on a comprehensive understanding of video and reference information. Furthermore, we propose a Language-Binded Reference Unification (LBRU) module to convert audio signals into language-formatted references, thereby unifying the formats of AVS and RVOS tasks in the same pipeline. Extensive experiments on both tasks show that our training-free AL-Ref-SAM 2 pipeline achieves performances comparable to or even better than fully-supervised fine-tuning methods. The code is available at: https://github.com/appletea233/AL-Ref-SAM2.
@@ Introduction
Figure 1: Comparison between the baseline and our method in training-free setting. (a) For the baseline, naively choosing the first frame to generate the object prompt may yield completely wrong predictions on the video since the first frame does not contain any relevant object. (b) Our method leverages GPT-4 to perform two-step temporal-spatial reasoning, selecting the frame and box that best reflects the reference information. The selected box serves as a more accurate object prompt to SAM 2 for better segmentation results.
@@
Referring Video Object Segmentation (RVOS) (Seo, Lee, and Han 2020) aims to segment the target object matched with the description of a given sentence. ReferFormer (Wu et al. 2022) and MTTR (Botach, Zheltonozhskii, and Baskin 2022) utilize language as queries which attend to the relevant visual regions through Transformer models. Temporal information is also investigated to align object motion with language expression by reasoning across multiple temporal scales (Han et al. 2023) or temporal interaction between the global referent token and object queries (Tang, Zheng, and Yang 2023). LoSh (Yuan et al. 2024) proposes to utilize both long and short text expressions to understand the appearance and motion cues of the target object. In this paper, we tackle the RVOS task in the training-free setting by leveraging the temporal-spatial reasoning capacity of GPT.
@@
Foundation models are large, pre-trained models that serve as a general-purpose base for a wide range of downstream tasks across different domains. GPT-4 (Achiam et al. 2023) has garnered significant attention for its impressive conversational and reasoning capabilities. In addition, other open-source vision-language models (such as CLIP (Radford et al. 2021), LLaVA (Liu et al. 2024b), OPT (Zhang et al. 2022), Stable Diffusion (Rombach et al. 2022), etc.) have also demonstrated excellent multimodal understanding and generation abilities. In terms of visual foundation models, the well-known SAM (Kirillov et al. 2023), trained on over a billion masks, demonstrates powerful zero-shot image segmentation capabilities based on various prompts. SAM-Track (Cheng et al. 2023) and SAM 2 (Ravi et al. 2024) further extend the fine-grained perception ability of SAM to videos, enabling general segmentation and tracking of objects within video sequences. In this paper, we design a novel two-step temporal-spatial reasoning flow to exploit the strong task execution capability of foundation models for training-free multimodal video object segmentation.
@@
The overall architecture of our Audio-Language-Referenced SAM 2 pipeline (AL-Ref-SAM 2) in illustrated in Figure 2. Given a video clip consisting of $T$ frames, and either a language description or an audio clip specifying the objects to be segmented (termed the referent), the goal of our AL-Ref-SAM 2  pipeline is to obtain the mask sequence $\mathbf{M}_{\rm{RVOS}}$ or $\mathbf{M}_{\rm{AVS}}\in\mathbb{R}^{T\times H\times W}$ of the referent across the whole video. Here, $H$ and $W$ denote the height and width of the video frames respectively. The language description is utilized in the RVOS task as the reference for indicating the referent. For the AVS task, we feed both the audio and video clips to the Language-Binded Reference Unification (LBRU) module to acquire a reference that describes the sounding objects from the acoustic aspect, thereby unifying the format of the AVS reference with that of the RVOS reference. Subsequently, the obtained reference and the video clip are processed by the GPT-assisted Pivot Selection (GPT-PS) module to identify a high-quality bounding box of the referent in a specific frame where the referent clearly appears. Finally, the selected bounding box serves as the pivot box to effectively prompt SAM 2 to segment the referent and propagate its mask forward and backward through the entire video clip. For the AVS task, since the selected referent may be silent in certain frames, we further utilize sound event detection to segment the whole audio clip and filter out silent frames from the obtained mask sequence, which is elaborated in the supplementary materials.
@@
Figure 3: Illustration of Language-Binded Reference Unification module.
@@
Step 2: Pivot Box Selection. As illustrated in the lower part of Figure 2(b), we first employ GroundingDINO to predict candidate boxes of the referent on the pivot frame according to the reference description. Since GroundingDINO is designed for grounding tasks on a single image, whereas our reference contains extensive temporal-related information, it is challenging to accurately predict the target object based solely on a single frame. Thus, we lower the confidence threshold for GroundingDINO to generate multiple candidates, maximizing the likelihood of the inclusion of the actual referent. Subsequently, we paint the candidate boxes on the pivot frame and mark each with a box ID to designate the regions for GPT’s attention. We also sequentially concatenate the other sampled frames with the marked pivot frame to provide temporal context for distinguishing the referent from multiple candidates. To further aid this process, a Pivot-Box CoT prompt, $\rm\mathbf{X}_{pb}$, is designed to guide GPT step-by-step in reasoning about the pivot box that best matches the reference description on the pivot frame. Specifically, based on the prior knowledge of the video event obtained in step 1, $\rm\mathbf{X}_{pb}$ first instructs GPT to describe the objects in each box, enabling it to comprehend candidates’ appearance, motion, and interrelationships. Afterward, GPT is required to perform syntactic analysis of the reference description to accurately identify the subject being referred to. Finally, the candidate box that best corresponds to the description of the identified subject is selected as the pivot box $p_{b}$. For illustrative purposes, we present an abridged version of the prompts in Figure 2(b). The complete version can be found in the supplementary materials.
@@
We adopt Ref-YouTube-VOS (Seo, Lee, and Han 2020), Ref-DAVIS17 (Khoreva, Rohrbach, and Schiele 2019), MeViS (Ding et al. 2023) for RVOS evaluation, and AVSBench (Zhou et al. 2022) datasets for AVS evaluation respectively. Details of these datasets are included in supplementary materials. In terms of evaluation metrics, we adopt region similarity $\mathcal{J}$ (average IoU), contour accuracy $\mathcal{F}$ and their average $\mathcal{J}\&\mathcal{F}$ for RVOS. The metrics $\mathcal{M_{J}}$ and $\mathcal{M_{F}}$ for AVS are the same as $\mathcal{J}$ and $\mathcal{F}$.
@@
In Figure 4, we compare the qualitative results of our method with those of the baseline (GroundingDINO + SAM 2) on the Ref-YouTube-VOS dataset. These results demonstrate that our method has a stronger ability to distinguish between objects of the same category compared to the baseline. For example, in the first row, even though only a small part of the bicycle’s handlebar is visible, our method can still accurately locate the referred bicycle in the back based on cues like “behind”. More qualitative results on the RVOS and AVS tasks are included in the supplementary materials.
@@ Appendix B Audio Clip Segmentation
In scenarios where multiple objects may produce sounds simultaneously or intermittently within the same video sample, it is crucial to determine the exact time intervals for each sounding object after obtaining its mask sequence throughout the entire video. To this end, we employ the combination of sound event detection (SED) and audio-language feature matching to associate each audio segment with one or more sounding object labels. Specifically, we first use a lightweight SED tool***https://github.com/robertanto/Real-Time-Sound-Event-Detection to split the audio $\bm{A}$ into $s$ segments $\{\bm{A}^{i}\}_{i=1}^{s}$ based on inter-frame feature similarity. Then, we extract the audio feature for each segment $\bm{f}_{a}^{i}\in\mathbb{R}^{C}$ using the audio encoder of LanguageBind (Zhu et al. 2023), where $C$ denotes the channel number of the audio features. Let $\{\bm{L}^{j}\}_{j=1}^{d}$ be the text labels derived from all $d$ possible combinations of object categories predicted by LBRU, we extract the text feature for each combination as $\bm{f}_{t}^{j}\in\mathbb{R}^{C}$ by feeding the textual labels to the text encoder of LanguageBind. Finally, since the audio features and the text features are spatially aligned, we compute the similarity between the audio features of each segment and the corresponding text features of the label combinations to associate the audio segment with the corresponding text label combination:

### v2

@@ Abstract
In this paper, we propose an Audio-Language-Referenced SAM 2 (AL-Ref-SAM 2) pipeline to explore the training-free paradigm for audio and language-referenced video object segmentation, namely AVS and RVOS tasks. The intuitive solution leverages GroundingDINO to identify the target object from a single frame and SAM 2 to segment the identified object throughout the video, which is less robust to spatiotemporal variations due to a lack of video context exploration. Thus, in our AL-Ref-SAM 2 pipeline, we propose a novel GPT-assisted Pivot Selection (GPT-PS) module to instruct GPT-4 to perform two-step temporal-spatial reasoning for sequentially selecting pivot frames and pivot boxes, thereby providing SAM 2 with a high-quality initial object prompt. Within GPT-PS, two task-specific Chain-of-Thought prompts are designed to unleash GPT’s temporal-spatial reasoning capacity by guiding GPT to make selections based on a comprehensive understanding of video and reference information. Furthermore, we propose a Language-Binded Reference Unification (LBRU) module to convert audio signals into language-formatted references, thereby unifying the formats of AVS and RVOS tasks in the same pipeline. Extensive experiments show that our training-free AL-Ref-SAM 2 pipeline achieves performances comparable to or even better than fully-supervised fine-tuning methods.
Codehttps://github.com/appletea233/AL-Ref-SAM2
@@ Introduction
Figure 1: (a) For the training-free baseline, naively choosing the first frame to generate the object prompt may yield completely wrong predictions on the video since the first frame does not contain any relevant object. (b) Our method leverages GPT-4 to perform two-step temporal-spatial reasoning, selecting the frame and box that best reflects the reference information. The selected box serves as a more accurate object prompt to SAM 2 for better segmentation results.
@@
Different from Referring Image Segmentation (Hui et al. 2020; Liu et al. 2021a; Huang et al. 2020; Liu, Jiang, and Ding 2024), Referring Video Object Segmentation (RVOS) (Seo, Lee, and Han 2020; Hui et al. 2021; Ding et al. 2022; Hui et al. 2023) aims to segment the target object matched with the description of a given sentence. ReferFormer (Wu et al. 2022) and MTTR (Botach, Zheltonozhskii, and Baskin 2022) utilize language as queries which attend to the relevant visual regions through Transformer models. Temporal information is also investigated to align object motion with language expression by reasoning across multiple temporal scales (Han et al. 2023; Tang et al. 2021) or temporal interaction between the global referent token and object queries (Tang, Zheng, and Yang 2023). LoSh (Yuan et al. 2024) proposes to utilize both long and short text expressions to understand the appearance and motion cues of the target object. In this paper, we tackle the RVOS task in the training-free setting by leveraging the temporal-spatial reasoning capacity of GPT.
@@
Foundation models are large, pre-trained models that serve as a general-purpose base for a wide range of downstream tasks across different domains (Shi et al. 2024; Liu et al. 2021b; Wang et al. 2023). GPT-4 (Achiam et al. 2023) has garnered significant attention for its impressive conversational and reasoning capabilities. In addition, other open-source vision-language models (such as CLIP (Radford et al. 2021), LLaVA (Liu et al. 2024b), OPT (Zhang et al. 2022), Diffusion Models (Rombach et al. 2022; Zhang et al. 2023), etc.) have also demonstrated excellent multimodal understanding and generation abilities. In terms of visual foundation models, the well-known SAM (Kirillov et al. 2023), trained on over a billion masks, demonstrates powerful zero-shot image segmentation capabilities based on various prompts. SAM-Track (Cheng et al. 2023) and SAM 2 (Ravi et al. 2024) further extend the fine-grained perception ability of SAM to videos, enabling general segmentation and tracking of objects within video sequences. In this paper, we design a novel two-step temporal-spatial reasoning flow to exploit the strong task execution capability of foundation models for training-free multimodal video object segmentation.
@@
The overall architecture of our Audio-Language-Referenced SAM 2 pipeline (AL-Ref-SAM 2) in illustrated in Figure 2. Given a video clip consisting of $T$ frames, and either a language description or an audio clip specifying the objects to be segmented (termed the referent), the goal of our AL-Ref-SAM 2  pipeline is to obtain the mask sequence $\mathbf{M}_{\rm{RVOS}}$ or $\mathbf{M}_{\rm{AVS}}\in\mathbb{R}^{T\times H\times W}$ of the referent across the whole video. Here, $H$ and $W$ denote the height and width of the video frames respectively. The language description is utilized in the RVOS task as the reference for indicating the referent. For the AVS task, we feed both the audio and video clips to the Language-Binded Reference Unification (LBRU) module to acquire a reference that describes the sounding objects from the acoustic aspect, thereby unifying the format of the AVS reference with that of the RVOS reference. Subsequently, the obtained reference and the video clip are processed by the GPT-assisted Pivot Selection (GPT-PS) module to identify a high-quality bounding box of the referent in a specific frame where the referent clearly appears. Finally, the selected bounding box serves as the pivot box to effectively prompt SAM 2 to segment the referent and propagate its mask forward and backward through the entire video clip. For the AVS task, we use sound event detection to segment the audio clip and filter out silent frames.
@@
Figure 3: Detailed illustration of our Language-Binded Reference Unification module.
@@
Step 2: Pivot Box Selection. As illustrated in the lower part of Figure 2(b), we first employ GroundingDINO to predict candidate boxes of the referent on the pivot frame according to the reference description. Since GroundingDINO is designed for grounding tasks on a single image, whereas our reference contains extensive temporal-related information, it is challenging to accurately predict the target object based solely on a single frame. Thus, we lower the confidence threshold for GroundingDINO to generate multiple candidates, maximizing the likelihood of the inclusion of the actual referent. Subsequently, we paint the candidate boxes on the pivot frame and mark each with a box ID to designate the regions for GPT’s attention. We also sequentially concatenate the other sampled frames with the marked pivot frame to provide temporal context for distinguishing the referent from multiple candidates. To further aid this process, a Pivot-Box CoT prompt, $\rm\mathbf{X}_{pb}$, is designed to guide GPT step-by-step in reasoning about the pivot box that best matches the reference description on the pivot frame. Specifically, based on the prior knowledge of the video event obtained in step 1, $\rm\mathbf{X}_{pb}$ first instructs GPT to describe the objects in each box, enabling it to comprehend candidates’ appearance, motion, and interrelationships. Afterward, GPT is required to perform syntactic analysis of the reference description to accurately identify the subject being referred to. Finally, the candidate box that best corresponds to the description of the identified subject is selected as the pivot box $p_{b}$. For illustrative purposes, we present an abridged version of the prompts in Figure 2(b).
@@
We adopt Ref-YouTube-VOS (Seo, Lee, and Han 2020), Ref-DAVIS17 (Khoreva, Rohrbach, and Schiele 2019), MeViS (Ding et al. 2023) for RVOS evaluation, and AVSBench (Zhou et al. 2022) datasets for AVS evaluation respectively. In terms of evaluation metrics, we adopt region similarity $\mathcal{J}$ (average IoU), contour accuracy $\mathcal{F}$ and their average $\mathcal{J}\&\mathcal{F}$ for RVOS. The metrics $\mathcal{M_{J}}$ and $\mathcal{M_{F}}$ for AVS are the same as $\mathcal{J}$ and $\mathcal{F}$.
@@
In Figure 4, we compare the qualitative results of our method with those of the baseline (GroundingDINO + SAM 2) on the Ref-YouTube-VOS dataset. These results demonstrate that our method has a stronger ability to distinguish between objects of the same category compared to the baseline. In the first row, even though only a small part of the bicycle’s handlebar is visible, our method still accurately locates the referred bicycle in the back based on cues like “behind”.
@@ Conclusion
Acknowledgments
This research was supported in part by National Science and Technology Major Project (2022ZD0115502), National Natural Science Foundation of China (No. U23B2010, No. 62461160308), Zhejiang Provincial Natural Science Foundation of China (Grant No. LDT23F02022F02), Beijing Natural Science Foundation (No. L231011), Beihang World TOP University Cooperation Program, and Meituan.
@@ Appendix B Audio Clip Segmentation
In scenarios where multiple objects may produce sounds simultaneously or intermittently within the same video sample, it is crucial to determine the exact time intervals for each sounding object after obtaining its mask sequence throughout the entire video. To this end, we employ the combination of sound event detection (SED) and audio-language feature matching to associate each audio segment with one or more sounding object labels. Specifically, we first use a lightweight SED tool [^1] https://github.com/robertanto/Real-Time-Sound-Event-Detection to split the audio $\bm{A}$ into $s$ segments $\{\bm{A}^{i}\}_{i=1}^{s}$ based on inter-frame feature similarity. Then, we extract the audio feature for each segment $\bm{f}_{a}^{i}\in\mathbb{R}^{C}$ using the audio encoder of LanguageBind (Zhu et al. 2023), where $C$ denotes the channel number of the audio features. Let $\{\bm{L}^{j}\}_{j=1}^{d}$ be the text labels derived from all $d$ possible combinations of object categories predicted by LBRU, we extract the text feature for each combination as $\bm{f}_{t}^{j}\in\mathbb{R}^{C}$ by feeding the textual labels to the text encoder of LanguageBind. Finally, since the audio features and the text features are spatially aligned, we compute the similarity between the audio features of each segment and the corresponding text features of the label combinations to associate the audio segment with the corresponding text label combination: